# DexCap: Scalable and Portable Mocap Data Collection System for Dexterous Manipulation

### v1

@@ Acknowledgments
This research was supported by National Science Foundation NSF-FRR-2153854 and Stanford Institute for Human-Centered Artificial Intelligence, SUHAI. This work is partially supported by ONR MURI N00014-21-1-2801. We would like to thank Yunfan Jiang, Albert Wu, Paul de La Sayette, Ruocheng Wang, Sirui Chen, Josiah Wong, Wenlong Huang, Yanjie Ze, Christopher Agia, Jingyun Yang and the SVL PAIR group for providing help and feedback. We also thank Zhenjia Xu, Cheng Chi, Yifeng Zhu for their suggestions in tuning the robot controller. We especially thank Kenneth Shaw, Ananye Agrawal, Deepak Pathak for open-sourcing the LEAP Hand.

### v2

@@ Acknowledgments
This research was supported by National Science Foundation NSF-FRR-2153854 and Stanford Institute for Human-Centered Artificial Intelligence, SUHAI. This work is partially supported by ONR MURI N00014-21-1-2801. We would like to thank Yunfan Jiang, Albert Wu, Paul de La Sayette, Ruocheng Wang, Sirui Chen, Josiah Wong, Wenlong Huang, Yanjie Ze, Christopher Agia, Jingyun Yang and the SVL PAIR group for providing help and feedback. We also thank Zhenjia Xu, Cheng Chi, Yifeng Zhu for their suggestions on the robot controller. We especially thank Kenneth Shaw, Ananye Agrawal, Deepak Pathak for open-sourcing the LEAP Hand.
@@ A-G Human-in-the-loop implementations
Appendix B Supplementary Experiment Results
Figure 15: Compare with IMU-based mocap system. We disable the SLAM mapping and pose-correction features of the T265 tracking camera, forcing it to rely on IMU information to track the pose. The human operator held the camera, started from a fixed location, moved it along a predefined trajectory, and then returned to the starting position. IMU-based method (first row) fails to match the endpoint with the start point, which indicates that there is pose drift during tracking. Our SLAM-IMU method (second row) doesn’t drift and captures smooth trajectory during the tracking.
Drifting error (cm) Trajectory 1 Trajectory 2
IMU-based $8.0\pm 3.1$ $11.3\pm 4.7$
SLAM-IMU (Ours) $0.4\pm 0.2$ $0.8\pm 0.3$
TABLE VIII: Drifting error of different tracking methods.
B-A Tracking accuracy
Figure 15 and Table VIII present qualitative and quantitative results, respectively. We observe that the IMU-based method suffers from pose drifting during tracking, while our SLAM-IMU approach more accurately tracks hand poses, with an average error of $0.8$ cm compared to the $11.3$ cm error of the IMU-based method.